# An Empirical Evaluation of Pre-trained Large Language Models for Repairing Declarative Formal Specifications

### v1

@@ Abstract.
Automatic Program Repair (APR) has garnered significant attention as a practical research domain focused on automatically fixing bugs in programs. While existing APR techniques primarily target imperative programming languages like C and Java, there is a growing need for effective solutions applicable to declarative software specification languages. This paper presents a systematic investigation into the capacity of Large Language Models (LLMs) for repairing declarative specifications in Alloy, a declarative formal language used for software specification. We propose a novel repair pipeline that integrates a dual-agent LLM framework, comprising a Repair Agent and a Prompt Agent. Through extensive empirical evaluation, we compare the effectiveness of LLM-based repair with state-of-the-art Alloy APR techniques on a comprehensive set of benchmarks. Our study reveals that LLMs, particularly GPT-4 variants, outperform existing techniques in terms of repair efficacy, albeit with a marginal increase in runtime and token usage. This research contributes to advancing the field of automatic repair for declarative specifications and highlights the promising potential of LLMs in this domain.
Declarative specification, Automatic Program Repair, formal methods, Alloy language, LLMs
copyright: none
1. Introduction
@@
Declarative specification languages have become instrumental in addressing a multitude of software engineering challenges. Among these languages, the Alloy specification language (Jackson, 2006) has emerged as a powerful tool, leveraging relational algebra and first-order logic to tackle a diverse array of tasks within the software engineering domain. Its application spans across software verification (Bagheri et al., [n. d.]), the security analysis of cutting-edge platforms such as the IoT and Android systems (Bagheri et al., 2016, 2021; Alhanahnah et al., [n. d.]), and the creation of test cases (Khurshid and Marinov, 2004; Mirzaei et al., 2016). The integration of Alloy with the Alloy Analyzer has facilitated automated property verification, simplifying the process of checking whether specifications adhere to desired properties and seamlessly integrating it into the Alloy environment. Despite these advancements, Alloy users, akin to developers in imperative languages, encounter challenges in debugging and correcting subtle bugs that may arise during specification writing, particularly for complex systems.
@@
While the Alloy Analyzer aids in automatic property verification and counterexample generation, debugging and rectifying issues in Alloy specifications remain laborious and manual tasks. Unlike the rich literature and techniques available for automatic program repair (APR) in imperative languages, the landscape for APR in declarative languages is relatively sparse. Early approaches like ARepair (Wang et al., 2019) primarily rely on failing test cases to identify and rectify bugs, assuming the presence of tests for verification. However, this paradigm does not align well with Alloy’s assertion-based specification approach, where users articulate expectations using assertions rather than tests. Existing APR techniques like ARepair (Wang et al., 2019) may succumb to overfitting issues when test coverage is insufficient, compromising the correctness of generated repairs. Despite efforts to mitigate these challenges with techniques like BeAFix (Brida et al., 2021), which leverages assertions as correctness oracles, there is room for improvement in terms of efficiency and effectiveness.
@@
The emergence of Large Language Models (LLMs) has revolutionized various domains, including natural language processing and code generation. These pre-trained models, such as GPT-3.5-Turbo and GPT-4, exhibit remarkable accuracy in predicting natural language and generating code. Inspired by these advancements and the persistent challenges in repairing formal specifications, we embark on a systematic empirical study to explore the potential of utilizing LLMs for repairing faulty Alloy specifications. Our research aims to address key questions regarding the effectiveness, performance, adaptive prompting, failure characteristics, and repair costs associated with employing LLMs in Alloy specification repair.
@@
This paper presents a methodical exploration and the results of our comprehensive evaluation of LLMs’ efficacy in repairing faulty Alloy specifications, juxtaposed against state-of-the-art Alloy APR techniques. We introduce a novel repair pipeline integrating a dual-agent LLM framework comprising a Repair Agent and a Prompt Agent  [^1] Our implementation and artifacts are available at: https://github.com/Mohannadcse/AlloySpecRepair. This iterative strategy centralizes around repairing bugs in defective specifications and generating specialized prompts to guide the repair process.
@@
We conduct an extensive evaluation of LLM effectiveness in repairing flawed Alloy specifications, benchmarking it against several state-of-the-art Alloy APR techniques, including ARepair(Wang et al., 2019), ICEBAR (Gutiérrez Brida et al., 2023), BeAFix (Brida et al., 2021), and ATR (Zheng et al., 2022), shedding light on their capabilities, benefits, and potential limitations. Our evaluation encompassed two comprehensive sets of benchmarks, comprising 1,974 defective specifications sourced from ARepair (Wang et al., 2019) and Alloy4Fun (Macedo et al., 2021), developed by external research groups. The repair pipeline facilitates the rectification of each defective specification under six distinct scenarios, with up to six iterations per scenario. Our findings contribute to advancing the field of automatic repair for declarative specifications and provide insights into leveraging LLMs for effective Alloy specification repair.
2. Background and Motivation
@@
This section provides an overview of multi-agent large language models followed by an illustrative example of a faulty Alloy specification to motivate the research and help the reader follow the discussions that ensue.
2.1. Multi-Agent Large Language Models
2.1.1. Large Language Models
@@
Emerging pre-trained Large Language Models (LLMs) have demonstrated impressive performance on natural language tasks, such as text generation (Brown et al., 2020; Chowdhery et al., 2022; Vaswani et al., 2017) and conversations (Thoppilan et al., 2022; OpenAI, 2023). LLMs have also been proven effective in translating natural language specifications and instructions into executable code (Fan et al., 2023; Jain et al., 2022; Che, 2021). These models have been trained on extensive corpora and possess the ability to execute specific tasks without the need for additional training or hard coding (Bubeck et al., 2023). They are invoked and controlled simply by providing a natural language prompt. The degree to which LLMs understand tasks depends largely on the prompting techniques used to convey user-provided instructions. These prompts are categorized into two frameworks: “zero-shot” learning and “few-shot” learning (Brown et al., 2020). Within the “zero-shot” learning framework, the prompt includes a description of the problem and the instructions for its resolution, aiming that LLMs can tackle previously unseen tasks. Conversely, in the context of “few-shot” learning, LLMs are provided with examples, supplementing the guidance offered in the “zero-shot” prompt.
2.1.2. Multi-Agent LLMs.
@@
With recent advancements in LLMs, developers are embracing the idea of creating autonomous agents that can solve various tasks and interact with environments, humans, and other agents using natural language interfaces (Zhou et al., 2023a). These agents provide several features including planning, memory, tool usage, and multi-agent communication. Therefore, LLM-based autonomous agents have gained tremendous interest in industry and academia (Wang et al., 2023). Several frameworks have been developed to support harnessing multi-agent LLM applications. AGENTS (Zhou et al., 2023b) is a unified framework and open-source library for language agents. AGENTS aims to facilitate developers to build applications with language agents. AutoGPT (Aut, [n. d.]) a multi-agent framework for LLMs, is designed to support autonomous applications of LLMs. LangChain (Chase, 2022) supports end-to-end language agents that can automatically solve tasks specified in natural language. It also facilitates the connection between LLM agents and external tools. Langroid (lan, 2024) is another framework that supports the development of multi-agent LLMs. It offers the capability to manage the history of messages, thus controlling the context window of LLMs. This is a crucial functionality for LLM apps operating iteratively (i.e., Self-Refine (Madaan et al., 2023) and Reflexion (Shinn et al., 2024)). Moreover, Langroid enables seamless integration with a variety of LLMs.
2.1.3. tools and function calling
@@
2.2. Motivating Example
Alloy, a formal modeling language, provides a comprehensible syntax inspired by object-oriented notations and is grounded on first-order relational logic (All, [n. d.]). Within an Alloy specification, three primary components shape its structure: data types, constraints expressed through formulas, and commands to initiate the analyzer.
The language relies on sig to define fundamental data types, which can be organized using fields to capture relationships between signatures. Additionally, Alloy employs fact to introduce constraints that persist in every instance of the specification. These constraints restrict the model space, ensuring its consistency. Further structuring of formulas is achieved through pred and fun, which are named parameterized Alloy formulas, and assert encapsulates the properties intended for validation.
@@ 2.2. Motivating Example
To illustrate the Alloy language and motivate the research, we examine the specification of the River Crossing puzzle, a classic problem of safely transporting a farmer, a fox, a chicken, and a sack of grain across a river. Listing 1 presents an abbreviated Alloy specification snippet of the farmer model from ARepair benchmark (Wang et al., 2019). In the River Crossing puzzle, the goal is to transport all objects from one side of the river to the other safely. However, there are constraints to consider: the farmer can only take one item with them in the boat, and if the farmer leaves the fox with the chicken, or the chicken with the grain, without supervision, a dire consequence will occur. This puzzle’s Alloy specification defines objects like the Farmer, Fox, Chicken, and Grain, along with their attributes and constraints.
However, the model contains a flaw in the crossRiver predicate, where the logic mistakenly assumes the occurrence of eating actions on the wrong side of the river. This oversight allows scenarios where the fox can eat the chicken without the farmer’s supervision, violating the puzzle’s rules. Interestingly, state-of-the-art automated repair techniques have struggled to address this particular bug. To repair this model, the erroneous logic within the crossRiver predicate must be corrected. However, this task presents a significant challenge due to the complexity of the puzzle’s logic and interactions between its components. Rectifying these logical errors requires a deep understanding of the problem domain and intricate relationships within the specification.
The objective of our study is to assess the effectiveness of repair methods for such buggy specifications using LLMs compared to state-of-the-art Alloy repair techniques. LLMs offer potential in uncovering complex violations and evaluating repair methods by reasoning about broader system contexts and intended behaviors. In the following sections, we explore the ability of LLMs to address such challenges and the specific conditions under which they can do so.
1abstract sig Object { eats: set Object }
2one sig Farmer, Fox, Chicken, Grain extends Object {}
3fact eating { eats = Fox->Chicken + Chicken->Grain }
4sig State {
5 near: set Object,
6 far: set Object
@@
8fact initialState {
9 let s0 = ord/first |
10 s0.near = Object && no s0.far
@@
12// Bug found in original model crossRiver
13pred crossRiver [from, from’, to, to’: set Object] {
14 // Fix: replace "from’ = from - Farmer && to’ = to - to.eats + Farmer" with "from’ = from - Farmer - from’.eats && to’ = to + Farmer".
15 (from = from - Farmer && to = to - to.eats + Farmer)
16 // Fix: replace "from’ = from - Farmer - item && to’ = to - to.eats + Farmer + item" with "from’ = from - Farmer - item - from’.eats && to’ = to + Farmer + item".
17 || (some item: from - Farmer {
18 from = from - Farmer - item
19 to = to - to.eats + Farmer + item })
@@
21fact stateTransition {
22 all s: State, s’: ord/next[s] {
23 Farmer in s.near =>
24 crossRiver[s.near, s’.near, s.far, s’.far] else
25 crossRiver[s.far, s’.far, s.near, s’.near]
26 }
@@
28pred solvePuzzle { ord/last.far = Object }
29run solvePuzzle for 8 State expect 1
30assert NoQuantumObjects {
31 no s : State | some x : Object | x in s.near and x in s.far
32}
33check NoQuantumObjects for 8 State expect 0
@@
Listing 1: A snippet of the flawed farmer Alloy specification sourced from the ARepair benchmark [^4] https://github.com/guolong-zheng/atmprep/blob/master/benchmark/arepair/farmer1.als.
3. Methodology
@@
This section outlines the repair pipeline, detailing its architecture and workflow. The pipeline is driven by the APR phases as outlined in (Zhang et al., 2023), including localization, repair, and verification. Our implementation of this pipeline utilizes the Langroid framework (lan, 2024).
Figure 1. LLM-assisted APR Pipeline.
3.1. Overall Workflow
@@
In this section, we overview our repair pipeline specifically tailored for iterative self-refinement. Our approach revolves around the concept of iterative refinement, wherein the repair process for a defective Alloy model operates within a predetermined budget, defined by the number of iterations allocated. Previous research has indicated that this iterative prompting method yields superior outcomes compared to traditional single-step approaches, enhancing task efficiency by an average of 20
@@
The workflow of our Automatic Program Repair (APR) pipeline, as depicted in Figure 1, comprises two primary components: the Repair Agent and the Prompt Agent. Each agent maintains its context, tools, and prompts, but they collaborate to adaptively refine the prompt for enhanced effectiveness, as elaborated below. Our pipeline supports various feedback levels, employing a zero-shot prompting approach (Paul et al., 2023), where tasks are performed without explicit examples, as detailed in Table 1.
@@
Upon receiving defective Alloy specifications and the prompt, our APR pipeline dispatches these elements to the LLM, which generates a revised version aimed at rectifying the defect. Subsequently, the Repair Agent compares the proposed specifications with the buggy ones to evaluate LLM’s ability to localize faults and adhere to instructions. The correctness of the proposed specifications is then assessed by triggering the Alloy Analyzer tool, which runs the proposed specifications for validation. If the Alloy analyzer confirms the bug’s resolution, the repair process concludes. Conversely, if errors persist, the pipeline initiates another iteration, contingent upon the remaining repair budget. If the budget is depleted, the pipeline terminates operations.
Our adaptive prompt is designed to evolve and refine based on different settings (detailed in Section 3.3), ensuring continual improvement after each unsuccessful attempt.
The following sections provide a comprehensive overview of each component of the pipeline.
Table 1. Zero-shot prompts used by the repair pipeline
@@
Repair Agent Agent-Instruction You are an expert in repairing Alloy declarative specifications. You will be presented with Alloy ¡Faulty_SPECIFICATIONS¿. Your task is to FIX/REPAIR the ¡Faulty_SPECIFICATIONS¿. Use the tool ‘run_alloy_analyzer‘ to demonstrate and validate the ¡FIXED_SPECIFICATIONS¿. Wait for my feedback, which may include error messages or Alloy solver results. You will have 5 trials to fix the ¡Faulty_SPECIFICATIONS¿. **Adhere to the Following Rules**: - The ¡FIXED_SPECIFICATIONS¿ should be consistent (having instances) and all the assertions should be valid (no counterexample). - DO NOT REPEAT the ¡FIXED_SPECIFICATIONS¿ that I sent you. - DO NOT provide any commentary and always send me anything ONLY using the tool ‘run_alloy_analyzer‘. - The ¡FIXED_SPECIFICATIONS¿ MUST be different than the ¡Faulty_SPECIFICATIONS¿.
Tool-instruction ALL AVAILABLE TOOLS and THEIR JSON FORMAT INSTRUCTIONS: You have access to the following TOOLS to accomplish your task: TOOL: run_alloy_analyzer, PURPOSE: To show a ¡FIXED_SPECIFICATIONS¿ to the user. Use this tool whenever you want to SHOW or VALIDATE the ¡FIXED_SPECIFICATIONS¿. NEVER list out a ¡FIXED_SPECIFICATIONS¿ without using this tool. JSON FORMAT: ”type”: ”object”, ”properties”: ”request”: ”default”: ”run_alloy_analyzer”, ”type”: ”string”,”specification”: ”type”: ”string”, ”required”: [ ”specification”, ”request”]. When one of the above TOOLs is applicable, you must express your request as ”TOOL:” followed by the request in the above JSON format.
Prompt Agent Agent-Instruction You are Expert in Analyzing Alloy Analyzer reports. Can you describe concisely and precisely the modifications needed to fix the error in at most 2 sentences? Based on this report from Alloy Analyzer: {Alloy_report_msg} After running this Alloy Model is: {proposed_spec}
3.2. Pipeline Agents

### v2

@@ Abstract
Automatic Program Repair (APR) has garnered significant attention as a practical research domain focused on automatically fixing bugs in programs. While existing APR techniques primarily target imperative programming languages like C and Java, there is a growing need for effective solutions applicable to declarative software specification languages. This paper systematically investigates the capacity of Large Language Models (LLMs) to repair declarative specifications in Alloy, a declarative formal language used for software specification. We designed 12 different repair settings, encompassing single-agent and dual-agent paradigms, utilizing various LLMs. These configurations also incorporate different levels of feedback, including an auto-prompting mechanism for generating prompts autonomously using LLMs. Our study reveals that dual-agent with auto-prompting setup outperforms the other settings, albeit with a marginal increase in the number of iterations and token usage. This dual-agent setup demonstrated superior effectiveness compared to state-of-the-art Alloy APR techniques when evaluated on a comprehensive set of benchmarks. This work is the first to empirically evaluate LLM capabilities to repair declarative specifications, while taking into account recent trending LLM concepts such as LLM-based agents, feedback, auto-prompting, and tools, thus paving the way for future agent-based techniques in software engineering.
1 Introduction
@@
Declarative specification languages have become instrumental in addressing a multitude of software engineering challenges. Among these languages, the Alloy specification language Jackson (2006) has emerged as a powerful tool, leveraging relational algebra and first-order logic to tackle a diverse array of tasks within the software engineering domain. Its application spans across software verification Bagheri et al. (2020), the security analysis of cutting-edge platforms such as the IoT and Android systems Bagheri et al. (2016, 2021); Alhanahnah et al. (2020), and the creation of test cases Khurshid and Marinov (2004); Mirzaei et al. (2016). Integrating Alloy with the Alloy Analyzer has facilitated automated property verification, simplifying the process of checking whether specifications adhere to desired properties, and seamlessly integrating it into the Alloy environment. Despite these advancements, Alloy users, similar to developers in imperative languages, encounter challenges in debugging and correcting subtle bugs that may arise during specification writing, particularly for complex systems.
@@
While the Alloy Analyzer aids in automatic property verification and counterexample generation, debugging and rectifying issues in Alloy specifications remain laborious and manual tasks. Unlike the rich literature and techniques available for automatic program repair (APR) in imperative languages, the landscape for APR in declarative languages is relatively sparse. Early approaches like ARepair Wang et al. (2019) rely mainly on failing test cases to identify and rectify bugs, assuming the presence of tests for verification. However, this paradigm does not align well with Alloy’s assertion-based specification approach, where users articulate expectations using assertions rather than tests. Existing APR techniques like ARepair Wang et al. (2019) may succumb to overfitting issues when the test coverage is insufficient, compromising the correctness of generated repairs. Despite efforts to mitigate these challenges with techniques like BeAFix Brida et al. (2021), which leverages assertions as correctness oracles, there is room for improvement in terms of efficiency and effectiveness.
@@
The emergence of Large Language Models (LLMs) has revolutionized various domains, including natural language processing and code generation. These pre-trained models, such as GPT-3.5-Turbo and GPT-4, exhibit remarkable accuracy in predicting natural language and generating code. Inspired by these advancements and the persistent challenges in repairing formal specifications, we embark on a systematic empirical study to explore the potential of using LLMs to repair faulty alloy specifications. Our research aims to investigate key questions regarding the effectiveness, performance, adaptive prompting, failure characteristics, and repair costs associated with the use of LLM in Alloy specification repair. This investigation leverages recent advancements, including LLM-agents, tools, and feedback mechanisms.
@@
This paper presents a systematic exploration and results of our comprehensive evaluation of the efficacy of LLMs in repairing faulty Alloy specifications, compared to state-of-the-art Alloy APR techniques. We design a repair pipeline working iteratively and integrating a dual-agent LLM framework comprising a Repair Agent and an Instructor Agent  [^1] Our implementation and artifacts are available at: https://github.com/Mohannadcse/AlloySpecRepair. This iterative strategy centralizes around repairing bugs in defective specifications and generating specialized prompts to guide the repair process.
@@
We conduct an extensive evaluation of LLM effectiveness in repairing defective Alloy specifications, comparing it against several state-of-the-art Alloy APR techniques, including ARepairWang et al. (2019), ICEBAR Gutiérrez Brida et al. (2023), BeAFix Brida et al. (2021), and ATR Zheng et al. (2022), shedding light on their capabilities, benefits, and potential limitations. Our evaluation encompassed two comprehensive sets of benchmarks, comprising 1,974 defective specifications sourced from ARepair Wang et al. (2019) and Alloy4Fun Macedo et al. (2021), developed by external research groups. For each defective specification, we assess the repair pipeline across six distinct settings, conducting up to six iterations per setting. Our findings contribute to advancing the field of automatic repair for declarative specifications and provide insights into leveraging LLMs for effective Alloy specification repair.
2 Background and Motivation
@@
This section provides an overview of multi-agent LLMs followed by an illustrative example of a faulty Alloy specification to motivate the research and help the reader follow the discussions that ensue.
2.1 Multi-Agent Large Language Models
2.1.1 Large Language Models
@@
Emerging pre-trained LLMs have demonstrated impressive performance on natural language tasks, such as text generation Brown et al. (2020); Chowdhery et al. (2022); Vaswani et al. (2017) and conversations Thoppilan et al. (2022); OpenAI (2023b). LLMs have also been proven effective in translating natural language specifications and instructions into executable code Fan et al. (2023); Jain et al. (2022); Che (2021). These models have been trained on extensive corpora and possess the ability to execute specific tasks without the need for additional training or hard coding Bubeck et al. (2023). They are invoked and controlled simply by providing a natural language prompt. The degree to which LLMs understand tasks depends largely on the prompting techniques used to convey user-provided instructions. These prompts are categorized into two frameworks: “zero-shot” learning and “few-shot” learning Brown et al. (2020). Within the “zero-shot” learning framework, the prompt includes a description of the problem and instructions for its resolution, aiming that LLMs can tackle previously unseen tasks. Conversely, in the context of “few-shot” learning, LLMs are provided with examples, supplementing the guidance offered in the “zero-shot” prompt.
2.1.2 Agentic LLMs
@@
With recent advancements in LLMs, developers are embracing the idea of creating autonomous agents that can solve various tasks and interact with environments, humans, and other agents using natural language interfaces Zhou et al. (2023b). These agents provide several features including planning, memory, tool usage, and multi-agent communication. Therefore, LLM-based autonomous agents have gained tremendous interest in industry and academia Wang et al. (2023). Several frameworks have been developed to support harnessing multi-agent LLM applications. AGENTS Zhou et al. (2023a) is a unified framework and open-source library for language agents. AGENTS aims to facilitate developers to build applications with language agents. AutoGPT AutoGPT (2024) a multi-agent framework for LLMs, is designed to support autonomous applications of LLMs. LangChain Chase (2022) supports end-to-end language agents that can automatically solve tasks specified in natural language. It also facilitates the connection between LLM agents and external tools. Langroid lan (2024) is another framework that supports the development of multi-agent LLMs. It offers the capability to manage the history of messages, thus controlling the context window of LLMs. This is a crucial functionality for LLM apps operating iteratively (i.e., Self-Refine Madaan et al. (2023) and Reflexion Shinn et al. (2024)). Moreover, Langroid enables seamless integration with a variety of LLMs.
2.1.3 Tools and function calling
@@
2.2 Alloy - Illustrative Example
Alloy, a formal modeling language, provides a comprehensible syntax inspired by object-oriented notations and is grounded on first-order relational logic Jackson (2006). Within an Alloy specification, three primary components shape its structure: data types, constraints expressed through formulas, and commands to initiate the analyzer.
The language uses signature (sig) definitions to define sets of elements, with fields specifying relationships between these sets. Additionally, Alloy employs fact to introduce constraints that hold in every instance of the specification. These constraints restrict the model space, ensuring its consistency. Further structuring of formulas is achieved through pred and fun, which are named parameterized Alloy expression, and assert encapsulates the properties intended for verification.
@@ 2.2 Alloy - Illustrative Example
To illustrate the Alloy language and provide motivation for this research, we examine the specification shown in Listing 1 from the ARepair benchmark Wang et al. (2019). This model represents a university context with students, professors, classes, and assignments. The specification defines relationships such as teaching assistants (students assigned as assistants), instructors (professors), and the association of assignments with both classes and students. The predicate PolicyAllowsGrading determines who can grade assignments, allowing only instructors or teaching assistants (TAs) of a class. However, a bug in the policy allows a student to grade their own assignment if they are TA for the same class. The fix is explicitly preventing students from grading their own assignments by adding the condition (at line 15). This condition ensures a person “s” who qualifies as a grader (either as a TA or an instructor) must not be among the students assigned that specific assignment. The assertion repair_assert_1 and predicate repair_pred_1 at the end formalize and check the requirement that no individual is allowed to grade an assignment assigned to themselves.
Reasoning about and addressing this defect through direct prompting of LLMs can be challenging, as demonstrated in Hasan et al. (2023). Additionally, conventional state-of-the-art Alloy repair techniques, such as ARepair Wang et al. (2019) and BeAFix Brida et al. (2021), also struggled to effectively address this defect. Given recent advances in LLM agents and prompt engineering techniques, this study aims to evaluate the capabilities of LLMs in light of these developments to repair such cases and evaluate their effectiveness in contrast to state-of-the-art Alloy repair techniques. In the following sections, we explore the ability of LLMs to address such challenges and the specific conditions under which they can do so.
1abstract sig Person {}
2sig Student extends Person {}
3sig Professor extends Person {}
4sig Class {
5 assistant_for: set Student, // as in : "is TA for"
6 instructor_of: one Professor
@@
8sig Assignment {
9 associated_with: one Class,
10 assigned_to: some Student
@@
12
13pred PolicyAllowsGrading(s: Person, a: Assignment) {
14 s in a.associated_with.assistant_for or s in a.associated_with.instructor_of
15 // Fix: add "s !in a.assigned_to".
16}
17
18assert repair_assert_1 {
19 all s : Person | all a: Assignment | PolicyAllowsGrading[s, a] implies not s in a.assigned_to
@@
21
22check repair_assert_1
23
24pred repair_pred_1 {
25 all s : Person | all a: Assignment | PolicyAllowsGrading[s, a] implies not s in a.assigned_to
26}
@@
28run repair_pred_1
@@
Listing 1: A snippet of the flawed grade Alloy specification sourced from the ARepair benchmark [^4] https://github.com/guolong-zheng/atmprep/blob/master/benchmark/arepair/grade1.als.
2.3 Underconstrained and Overconstrained Specifications
2.3.1 Underconstrained Specifications
Underconstrained specifications arise when essential constraints are omitted, permitting behaviors that deviate from the intended semantics. This class of defects often results in models that allow for unintended or invalid configurations. As illustrated in Listing 2, the specification enforces that each Tree must have at least one root, but does not restrict the number of root nodes. Consequently, multiple nodes could simultaneously serve as roots for a single tree, contradicting the common assumption of a unique root in tree structures. Repairs addressing this issue typically involve replacing some with one, or introducing explicit uniqueness constraints to ensure the intended behavior.
1sig Tree {
2root: set Node
3}
4sig Node {
5children: set Node
6}
7fact WellFormed {
8all t: Tree | some t.root
9}
Listing 2: Example of an underconstrained specification that permits multiple roots.
2.3.2 Overconstrained Specifications
In contrast, overconstrained specifications impose overly restrictive conditions, unintentionally excluding valid model instances. This type of defect typically results in models that are too limited, preventing the representation of legitimate scenarios. As shown in Listing 3, while the first constraint correctly enforces a unique root, the second constraint forbids any Node from having children—effectively eliminating all hierarchical structures. As a result, the specification precludes valid tree configurations. Repairs in such cases often involve relaxing the restrictive constraint, such as substituting no with some, to reintroduce permissible structure while preserving correctness.
1fact WellFormed {
2all t: Tree | one t.root
3all n: Node | no n.children
4}
Listing 3: Example of an overconstrained specification that disallows any node hierarchy.
3 Methodology
@@
This section outlines the repair pipeline, detailing its architecture and workflow. The pipeline is derived from the APR phases described in Zhang et al. (2023), including localization, repair, and verification. The LLM is responsible for the localization and repair steps, while the verification step is facilitated by granting the LLM access to the Alloy analyzer. Consequently, the LLM can autonomously execute all APR phases to repair the defective model. Our implementation of this pipeline utilizes the Langroid framework lan (2024).
3.1 Overall Workflow
@@
This section presents the design of the repair pipeline specifically designed for iterative self-refinement, wherein the repair process for a defective Alloy model operates within a predetermined budget, defined by the number of iterations allocated. Previous research has indicated that this iterative prompting method produces superior results compared to traditional single-step approaches, improving task efficiency by an average of 20% Madaan et al. (2023). Should the repair process exceed this limit without success, the pipeline halts further attempts. After each unsuccessful repair iteration, feedback is collected and used to update the prompt, refining the initial draft generated by the LLM.
Figure 1: APR Pipeline for Alloy specifications. It supports Single-Agent (Repair Agent Only) and Dual-Agent (Repair Agent and Instructor Agent) paradigms.
@@
The workflow of our APR pipeline, as depicted in Figure 1, comprises two primary components: the Repair Agent and the Instructor Agent. Each agent maintains its context, tools, and prompts, but they collaborate to adaptively refine the Repair Agent prompt for enhanced effectiveness, as elaborated below. Our pipeline supports various feedback levels, employing a zero-shot prompting approach Paul et al. (2023), where tasks are performed without explicit examples, as detailed in Table 1.
@@
Upon receiving defective Alloy specifications and the prompt, our APR pipeline dispatches these elements to the LLM, which generates a patched version aimed at rectifying the defect. Subsequently, the Repair Agent compares the proposed model with the buggy ones to evaluate the ability of LLM to locate faults and adhere to instructions. The accuracy of the proposed model is then assessed by triggering the Alloy Analyzer tool, which runs the proposed model for validation. If the Alloy analyzer confirms the bug’s resolution, the repair process concludes. Conversely, if errors persist, the pipeline initiates another iteration, depending on the remaining repair budget. If the budget is depleted, the pipeline terminates operations. The adaptive prompt is designed to evolve and refine based on different settings (detailed in Section 3.3), ensuring continuous improvement after each unsuccessful attempt. The following sections provide a comprehensive overview of each component of the repair pipeline.
Table 1: Zero-shot prompts used by the repair pipeline
@@
Repair Agent Agent-Instruction You are an expert in repairing Alloy declarative specifications. You will be presented with Alloy $<$Faulty_SPECIFICATIONS$>$. Your task is to FIX/REPAIR the $<$Faulty_SPECIFICATIONS$>$. Use the tool ‘run_alloy_analyzer‘ to demonstrate and validate the $<$FIXED_SPECIFICATIONS$>$. Wait for my feedback, which may include error messages or Alloy solver results. You will have 5 trials to fix the $<$Faulty_SPECIFICATIONS$>$. **Adhere to the Following Rules**: - The $<$FIXED_SPECIFICATIONS$>$ should be consistent (having instances) and all the assertions should be valid (no counterexample). - DO NOT REPEAT the $<$FIXED_SPECIFICATIONS$>$ that I sent you. - DO NOT provide any commentary and always send me anything ONLY using the tool ‘run_alloy_analyzer‘. - The $<$FIXED_SPECIFICATIONS$>$ MUST be different than the $<$Faulty_SPECIFICATIONS$>$.
Tool-instruction ALL AVAILABLE TOOLS and THEIR JSON FORMAT INSTRUCTIONS: You have access to the following TOOLS to accomplish your task: TOOL: run_alloy_analyzer, PURPOSE: To show a $<$FIXED_SPECIFICATIONS$>$ to the user. Use this tool whenever you want to SHOW or VALIDATE the $<$FIXED_SPECIFICATIONS$>$. NEVER list out a $<$FIXED_SPECIFICATIONS$>$ without using this tool. JSON FORMAT: ”type”: ”object”, ”properties”: ”request”: ”default”: ”run_alloy_analyzer”, ”type”: ”string”,”specification”: ”type”: ”string”, ”required”: [ ”specification”, ”request”]. When one of the above TOOLs is applicable, you must express your request as ”TOOL:” followed by the request in the above JSON format.
Instructor Agent Agent-Instruction You are Expert in Analyzing Alloy Analyzer reports. Can you describe concisely and precisely the modifications needed to fix the error in at most 2 sentences? Based on this report from Alloy Analyzer: {Alloy_report_msg} After running this Alloy Model is: {proposed_spec}
3.2 Pipeline Agents